OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

Do u know: parallel lines have equal slopes

OpenStudy (anonymous):

@soccerstar808 ???

OpenStudy (anonymous):

yes but like this is confusing me to the point where i am gonna cry

OpenStudy (anonymous):

okay start there both lines have equal solpes, therefore we know that the gradient m=4

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

then we know that the eqn of a straight line is y=mx + c

OpenStudy (anonymous):

so it would be 6=(4)-12+2?

OpenStudy (anonymous):

So can I help?

OpenStudy (anonymous):

sure any way you can i need the answer asap

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

x-4y=-18 but it doesnt seem rite

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

what is it then

OpenStudy (anonymous):

let y=4x+c be the equation of the required line then use the given point to find the value of c

OpenStudy (anonymous):

c=18

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

use the eqn y-y1=m(x-x1)

OpenStudy (anonymous):

so, y=4x+18

OpenStudy (anonymous):

is the correct answer

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh ok but the homework answers are x– 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18

OpenStudy (anonymous):

substitute given point in the question and then your answer pops out, two simple steps.

OpenStudy (anonymous):

y=4x+18 -18=4x-y 4x-y=-18

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep thanks.